# Disentangling Core and Edge Mechanisms of the Density Limit in DIII-D Negative Triangularity Plasmas

### v3

@@ Abstract
The density limit is investigated in the DIII-D negative triangularity (NT) plasmas which lack a standard H-mode edge. We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms. Sustained, non-disruptive operation is achieved at densities up to 1.8 times the Greenwald limit () until the termination of auxiliary heating. Systematic power scans reveal distinct power scalings for the core () and edge () density limits. The edge density saturation is triggered abruptly by the onset of a non-disruptive, high-field side radiative instability that clamps the edge density below . In contrast, the core density continues to rise until it saturates, a state characterized by substantially enhanced core turbulence. Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis ), elevated Hurst exponents, and a -type power spectrum. These findings suggest that the density limit in the low-confinement regime is determined by a combination of edge radiative instabilities and core turbulent transport. This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
@@ I Introduction
The conventional metric for the density limit is the empirical Greenwald scaling [2, 3]: , where denotes the plasma current and the minor radius. This scaling is notably independent of heating power. This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power. Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that the achievable density limit in the low-confinement (L-mode) regime may increase with heating power, potentially exceeding the conventional Greenwald density. Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physical basis rooted in power balance and transport that transcends specific magnetic configurations. Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
@@
The physics underpinning the L-mode density limit involves a complex interplay of several processes. Turbulent transport [11, 12, 13, 14, 15], radiative power losses [16, 7, 4], and magnetohydrodynamic (MHD) instabilities [17, 18] have all been reported as key contributors. These processes do not act in isolation; instead, they are coupled synergistically [3] and ultimately constrain the plasma density. Consequently, developing predictive models of density limits requires disentangling their individual roles and interdependencies.
@@
To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters. This avoids the complexities introduced by the edge transport barriers (ETBs) and edge localized modes (ELMs), which are characteristic of conventional high-confinement (H-mode) plasmas. The negative triangularity (NT) plasmas [19, 20], characterized by an inverted “D”-shaped cross-section (see Fig. 1), provides an ideal platform for this study. These plasmas exhibit improved core confinement [21, 22, 23, 24] while maintaining L-mode-like edge conditions. While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
@@
This paper presents an examination of the physics governing the density limit in NT plasmas on the DIII-D tokamak. We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power. A systematic power scan reveals disparate scaling dependencies for the core () and the separatrix () regions. Detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas. At the plasma edge, the density is constrained by enhanced turbulence driven by a radiative instability. In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence. These findings challenge the traditional paradigm of a single disruptive limit, advancing a more nuanced picture where distinct processes govern density saturation in different plasma regions.
@@ II Experimental Arrangement
The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power provided by neutral beam injection (NBI) and varied between discharges. All discharges employed a lower single-null configuration with an open divertor geometry, and the ion drift was directed away from the primary divertor, as illustrated in Fig. 1. The plasma shape was characterized by negative triangularity at both the top and bottom ( and ), moderate elongation (), and an edge safety factor of . This paper focuses specifically on discharges with toroidal field T and plasma current MA. Additional experiments exploring different plasma current and toroidal field combinations are reported separately [26]. In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber. Consistent with the behavior of NT plasmas, no edge localized modes were observed in any discharge.
A comprehensive suite of diagnostics monitored the evolving plasma state. Figure 1 shows the locations of the key diagnostics. Electron density and temperature in the main chamber were measured using the Thomson scattering (TS) system [27]. Charge exchange recombination (CER) spectroscopy [28] measured the temperature, density, and toroidal and poloidal rotation velocities of carbon ions. These impurity ion profiles were used to compute radial electric field profiles () using the carbon ion force balance equation. Bolometer arrays [29] provided measurements of the radiated power distribution and total radiated power.
Several diagnostics were used to probe density fluctuations. A multichannel interferometer [30] measured the line-averaged density and its low- fluctuations along one horizontal midplane chord and three vertical chords at major radii of (high-field side), (core region), and (low-field side). The radial interferometer polarimeter (RIP) [31] also provided line-integrated measurements of electron density fluctuations at the mid-plane. A 2D array of high-sensitivity beam emission spectroscopy (BES) [32] measured ion-scale () density fluctuations at the edge. Two 8-channel Doppler back-scattering (DBS) systems [33, 34] measured turbulent flows and low-to-intermediate- () density fluctuations. The radial locations and corresponding wavenumbers of the DBS measurements were calculated using the 3D ray-tracing code GENRAY [35].
@@ III.1.1 Density Saturation Beyond the Greenwald Limit
A key result of this study is the achievement of sustained L-mode operation at densities that surpass the conventional Greenwald density limit. This is shown in Fig. 2, which presents the time evolution of key parameters in two representative discharges with different levels of auxiliary heating from NBI. Across the experiment, the NBI power was varied from 3 MW up to 13 MW. As shown in Figs. 2(b) and 2(c), both core and separatrix densities increase with input power. The core density at , which is similar to the line-averaged density, consistently exceeds the Greenwald density (). Notably, rather than disrupting upon reaching their maximum density, the discharges with medium to high NBI power () transitioned into a phase of density saturation that persisted until the termination of auxiliary heating.
@@ III.1.2 Disparate Power Scaling for Core and Edge Density Limit
A systematic power scan reveals that the core and edge densities respond differently to , as shown in Fig. 3. The maximum core density (at ) exhibits a weak but clear power dependence, scaling as . This result is consistent with L-mode density limit scalings reported on other tokamaks, such as ASDEX Upgrade [8]. In contrast, the separatrix density displays a stronger relationship, scaling as . This relatively strong power dependence is reminiscent of the density limit observed in stellarators, where power balance is known to be the dominant constraint [9, 10].
These disparate power scalings constitute a central finding of this work; they indicate that the density saturation is not governed by a single, global mechanism. Instead, it suggests that distinct physical processes, with different sensitivities to heating power, are responsible for limiting the density in the core and edge regions. This hypothesis—that core saturation is dominated by turbulent transport while edge saturation is governed by radiative instabilities—guides the investigative framework for the remainder of this paper.
@@ III.2.1 Reduced Temperature and Rotational Shear at High Density
As the plasma approaches the density limit, its equilibrium profiles evolve in a manner that signals a clear degradation of thermal confinement. This evolution, shown for a representative high-power discharge in Fig. 4, provides important insight into the underlying transport dynamics.
The electron density profile [Fig. 4(a)] rises across the radius but remains notably flat in the core, indicating that strong particle transport persists even at high densities. The degraded confinement is accompanied by a loss of thermal energy, as both the electron and ion temperature profiles collapse [Figs. 4(b), 4(c)]. A noticeable feature is the pronounced flattening of the edge profile, a classic signature of intensified radiative cooling at the plasma boundary.
Another critical change occurs in the plasma rotation. The toroidal rotation profile [Fig. 4(d)], initially peaked in the core, progressively collapses with increasing density. This loss of core rotation directly dismantles the dominant component of the radial electric field (), a key factor for turbulence suppression. As shown in Fig. 5, the core shear, initially strong, diminishes and ultimately vanishes as the density limit is approached.
This cascade of events, from intensified edge cooling to the collapse of core rotation and shear, indicates that the plasma’s ability to suppress turbulence deteriorates with rising density. The loss of this stabilizing flow shear enables enhanced turbulent transport that governs the core density limit, as subsequent sections will demonstrate.
@@ III.2.2 Divergent Evolution of Turbulence in Different Regions
The degradation of thermal confinement described above coincides with a dramatic divergence in turbulent activity across the plasma, as illustrated in Fig. 6 for a high-power discharge. A critical transition occurs at approximately 3700 ms (vertical dashed line), marking the onset of strong edge cooling. Before this event, the turbulence evolution is relatively uniform; afterward, it diverges sharply between different spatial regions.
Initially, both core and edge densities rise together [Fig. 6(a)], accompanied by a steady increase in the amplitude of local, intermediate- turbulence at the low-field side (LFS) midplane (), as measured by Doppler backscattering [Fig. 6(d)]. This period of evolution ends abruptly at the 3700 ms transition. Immediately following this transition, the LFS turbulence saturates. This saturation coincides with the clamping of the edge density and the sharp drop in local electron temperature [Fig. 6(b)].
In contrast, turbulence in other regions intensifies. Line-integrated, low- density fluctuations on the high-field side (HFS), measured by the CO2 interferometer, grows substantially [Figs. 6(e), 6(f)]. A similar growth is observed in the line-integrated midplane fluctuations measured by the RIP diagnostic [Figs. 6(g), 6(h)].
This divergent evolution—saturation of localized turbulence on the LFS versus continued, strong growth of HFS and core fluctuations—reveals a profound poloidal asymmetry in the transport dynamics. This asymmetry strongly suggests that the density limit is not a globally uniform process but is instead triggered by a localized instability on the high-field side—likely the radiative phenomenon that we investigate next.
III.3 The Edge Density Limit: A Radiatively Driven Phenomenon
@@ III.3.1 Visualizing the HFS Radiation Front
The poloidally asymmetric turbulence described previously points to a localized radiative instability as the trigger for the edge density limit. Here, we examine the 2D radiated power distribution, reconstructed from bolometer measurements, to visualize the formation and evolution of this instability. Figure 7 presents a time-sequence of this evolution. While geometrically similar to a disruptive MARFE [16, 36], the feature observed here is a quasi-stable, non-disruptive radiation front that is directly linked to the observed density saturation.
The evolution of this front proceeds in clear stages. Early in the discharge ( ms), before the density ramp, the radiation is concentrated at the inner divertor target, characteristic of a standard attached divertor [Fig. 7(a)]. As plasma density increases ( ms), an intense radiation zone emerges near the X-point, signaling the birth of the radiative front [Fig. 7(b)]. This event coincides with the onset of the divergent turbulence seen in Fig. 6. With a further increase in density ( ms), this front detaches from the divertor and propagates upward along the HFS separatrix [Fig. 7(c)]. This movement is strongly correlated with the localized edge cooling and turbulence saturation. Finally ( ms), the front moves further inboard and its emissivity intensifies [Fig. 7(d)].
This sequence provides direct visual evidence of a developing radiative instability on the HFS. This dynamic structure appears to be the primary driver of the edge density limit. Notably, the eventual discharge termination was caused by a subsequent drop in NBI power that triggered an locked mode and was not a direct consequence of the radiation front.
@@ III.3.2 Radiative Cooling and the Onset of HFS Turbulence
The formation of the HFS radiation front actively cools the plasma and triggers a dramatic shift in the local turbulent state. Figure 8 provides quantitative evidence of this transition. As the HFS radiated power increases beyond a critical threshold, the edge plasma parameters and HFS turbulence changes fundamentally. The RMS amplitude of the line-integrated HFS density fluctuations (), measured by the CO2 interferometer passing through the radiation front, serves as a proxy for the instability’s strength.
The effect of edge cooling can be characterized by the dimensionless adiabaticity parameter [37], , which represents the comparison between parallel electron motion and turbulent timescales. Here, is the parallel wavenumber of turbulence, is the electron thermal velocity, is the turbulence frequency, and is the electron-ion collision frequency. When , the plasma enters a non-adiabatic regime where drift-wave instabilities are expected to grow. As shown in Fig. 8(a), the onset of the strong HFS radiation drives to plummet below 1.
This transition into a non-adiabatic state coincides with a surge in HFS turbulence. Figure 8(b) shows that the RMS amplitude of HFS density fluctuations () increases sharply as the radiated power crosses the same threshold. The entire event is correlated with a degradation in plasma performance, as indicated by the drop in the confinement factor (color scale). This finding demonstrates that the HFS radiation front alters the local plasma state, unleashing the turbulence that may ultimately limit the edge density.

### v4

@@ Abstract
The density limit is investigated in the DIII-D negative triangularity (NT) plasmas which lack a standard H-mode edge. We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms. Sustained, non-disruptive operation is achieved at densities up to 1.8 times the Greenwald limit () until the termination of auxiliary heating. Systematic power scans show distinct power scalings for the core () and edge () density limits. The edge density saturation is triggered by the onset of a non-disruptive, high-field side radiation front and the associated cooling, which clamps the edge density below . In contrast, the core density continues to rise until it saturates, a state characterized by enhanced core turbulence. Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis ), increased Hurst exponents, and a -type power spectrum. These findings suggest that the density limit in the low-confinement regime is determined by a combination of edge radiative cooling and core turbulent transport. This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
@@ I Introduction
The conventional metric for the density limit is the empirical Greenwald scaling [2, 3]: , where denotes the plasma current and the minor radius. This scaling is notably independent of heating power. This paradigm has been challenged by observations that the achievable density scales positively with heating power. Recent theoretical modeling [5, 6, 7] and database analyses [8, 9, 10] suggest that the achievable density limit in the low-confinement (L-mode) regime may increase with heating power, potentially exceeding the conventional Greenwald density. Furthermore, the power-dependent density limits have also been observed in stellarators [11, 12], which lack plasma current, suggesting a more universal physical basis rooted in power balance and transport that transcends specific magnetic configurations. Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
@@
The physics underpinning the L-mode density limit involves a complex interplay of several processes. Turbulent transport [13, 14, 15, 16, 17], radiative power losses [18, 7, 4], and magnetohydrodynamic (MHD) instabilities [19, 20] have all been reported as key contributors. These processes do not act in isolation; instead, they are coupled [3] and ultimately constrain the plasma density. Consequently, developing predictive models of density limits requires disentangling their individual roles and interdependencies.
@@
To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters. This avoids the complexities introduced by the edge transport barriers (ETBs) and edge localized modes (ELMs), which are characteristic of conventional high-confinement (H-mode) plasmas. The negative triangularity (NT) plasmas [21, 22], characterized by an inverted “D”-shaped cross-section (see Fig. 1), provides an ideal platform for this study. These plasmas exhibit improved core confinement [23, 24, 25, 26] while maintaining L-mode-like edge conditions. While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of H-mode ETBs and ELMs makes it a suitable proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
@@
This paper presents an examination of the physics governing the density limit in NT plasmas on the DIII-D tokamak. We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power. A systematic power scan reveals disparate scaling dependencies for the core () and the separatrix () regions. Detailed experimental measurements indicate that different mechanisms determine density limits in the core and edge plasmas. At the plasma edge, the density is constrained by enhanced turbulence driven by radiative cooling. In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence. These findings challenge the traditional paradigm of the density limit as a single boundary often associated with the empirical Greenwald scaling, advancing a more nuanced picture where distinct processes govern density saturation in different plasma regions.
In this work, the term ‘density limit’ refers to a multifaceted density saturation phenomenon—the point at which the plasma density ceases to rise despite continued gas puffing. This distinguishes our focus from studies where the ‘limit’ is defined by the onset of a disruption [10].
@@ II Experimental Arrangement
The experiment was conducted on the DIII-D tokamak [27], with auxiliary heating power provided by neutral beam injection (NBI) and varied between discharges. All discharges employed a lower single-null configuration with an open divertor geometry, and the ion drift was directed away from the primary divertor, as illustrated in Fig. 1. The plasma shape was characterized by negative triangularity at both the top and bottom ( and ), elongation (), and an edge safety factor of . This paper focuses specifically on discharges with toroidal field T and plasma current MA. Additional experiments exploring different plasma current and toroidal field combinations are reported in an accompanying paper [28]. In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber. Consistent with the behavior of NT plasmas, no edge localized modes were observed in any discharge.
A suite of diagnostics monitored the evolving plasma state. Figure 1 shows the locations of the key diagnostics. Electron density and temperature in the main chamber were measured using the Thomson scattering (TS) system [29]. Charge exchange recombination (CER) spectroscopy [30] measured the temperature, density, and toroidal and poloidal rotation velocities of carbon ions. These impurity ion profiles were used to compute radial electric field profiles () using the carbon ion force balance equation. Bolometer arrays [31] provided measurements of the radiated power distribution and total radiated power.
Several diagnostics were used to probe density fluctuations. A multichannel interferometer [32] measured the line-averaged density and its low- fluctuations along one horizontal midplane chord and three vertical chords at major radii of (high-field side), (core region), and (low-field side). The radial interferometer polarimeter (RIP) [33] also provided line-integrated measurements of electron density fluctuations at the mid-plane. A 2D array of beam emission spectroscopy (BES) [34] measured ion-scale () density fluctuations at the edge. Two 8-channel Doppler back-scattering (DBS) systems [35, 36] measured turbulent flows and low-to-intermediate- () density fluctuations. The radial locations and corresponding wavenumbers of the DBS measurements were calculated using the 3D ray-tracing code GENRAY [37].
@@ III.1.1 Density Saturation Beyond the Greenwald Limit
This study shows that the density limit in these plasmas can surpass the conventional Greenwald limit and manifests as a saturation phenomenon. This is shown in Fig. 2, which presents the time evolution of key parameters in two representative discharges with different levels of auxiliary heating from NBI. Across the experiment, the NBI power was varied from 3 MW up to 13 MW. As shown in Figs. 2(b) and 2(c), both core and separatrix densities increase with input power. The core density at , which serves as a proxy for the line-averaged density, exceeds the Greenwald density (). The discharges with medium to high NBI power () transitioned into a phase of density saturation that persisted until the termination of auxiliary heating.
@@ III.1.2 Disparate Power Scaling for Core and Edge Density Limit
A systematic power scan reveals that the core and edge densities respond differently to , as shown in Fig. 3. The maximum core density (at ) exhibits a power scaling of . In contrast, the separatrix density displays a stronger power dependence, scaling as . This is consistent with edge density limit scalings reported on other tokamaks; for example, it is in agreement with the scaling of recently reported by ASDEX Upgrade [8]. This relatively strong power dependence of the edge density limit is also reminiscent of the density limit observed in stellarators, where power balance is known to be the dominant constraint [11, 12].
These different power scalings indicate that the density saturation is not governed by a single, global mechanism. Instead, it suggests that distinct physical processes, with different sensitivities to heating power, are responsible for limiting the density in the core and edge regions.
@@ III.2.1 Reduced Temperature and Rotational Shear at High Density
As the plasma approaches the density limit, its equilibrium profiles evolve with degradation of thermal confinement. This evolution, shown for a representative high-power discharge in Fig. 4, provides further insight into the underlying transport dynamics.
The electron density profile [Fig. 4(a)] rises across the radius but remains flat in the core, indicating persistent particle transport at high densities. The degraded confinement is accompanied by a loss of thermal energy, as both the electron and ion temperature profiles collapse [Figs. 4(b), 4(c)]. The edge profile flattens significantly, a characteristic signature of intensified radiative cooling at the plasma boundary.
Another critical change occurs in the plasma rotation. The toroidal rotation profile [Fig. 4(d)], initially peaked in the core, progressively collapses with increasing density. This loss of core rotation eliminates the dominant component of the radial electric field (), reducing turbulence suppression of flow shear decorrelation. As shown in Fig. 5, the core shear, initially strong, diminishes and ultimately vanishes as the density limit is approached.
This sequence of events, from intensified edge cooling to the collapse of core rotation and shear, indicates that the plasma’s ability to suppress turbulence deteriorates with rising density. The loss of this stabilizing flow shear enables enhanced turbulent transport that governs the core density limit, as subsequent sections will demonstrate.
@@ III.2.2 Divergent Evolution of Turbulence in Different Regions
The degradation of thermal confinement described above coincides with divergence in turbulent activity across the plasma, as illustrated in Fig. 6 for a high-power discharge. A critical transition occurs at approximately 3750 ms (vertical dashed line), marking the onset of strong edge cooling. Before this event, the turbulence evolution is relatively uniform; afterward, it diverges sharply between different spatial regions.
Initially, both core and edge densities rise together [Fig. 6(a)], accompanied by a steady increase in the amplitude of local, intermediate- turbulence at the low-field side (LFS) midplane (), as measured by Doppler backscattering [Fig. 6(d)]. This period of evolution ends abruptly, initiated by a transition at approximately 3750 ms. Immediately following this trigger, the LFS turbulence saturates. A clamping of the edge density and a sharp drop in local electron temperature also start, which are fully established by ms [Fig. 6(b)].
In contrast, turbulence in other regions intensifies. Line-integrated, low- density fluctuations on the high-field side (HFS), measured by the CO2 interferometer, increase substantially [Figs. 6(e), 6(f)]. A similar growth is observed in the line-integrated midplane fluctuations measured by the RIP diagnostic [Figs. 6(g), 6(h)].
This divergent evolution—saturation of localized turbulence on the LFS versus continued, strong growth of HFS and core fluctuations—indicates a strong poloidal asymmetry in the transport dynamics. This asymmetry indicates that the density limit is not a uniform process but is likely triggered by a localized mode on the high-field side—likely the radiative phenomenon that we investigate next.
III.3 The Edge Density Saturation and the Role of the HFS Radiation Front
@@ III.3.1 Visualizing the HFS Radiation Front
The poloidally asymmetric turbulence described previously points to a localized HFS mode as the trigger for the edge density saturation. Here, we examine the 2D radiated power distribution, reconstructed from bolometer measurements, to visualize the formation and evolution of this mode. Figure 7 presents a time-sequence of this evolution. While geometrically similar to a disruptive MARFE [18, 38], the feature observed here is a quasi-stable, non-disruptive radiation front that is directly linked to the observed density saturation. Its onset is sometimes referred to as a radiative mode to emphasize its role in triggering the transition to a high-transport state, though it evolves into a persistent radiative structure reminiscent of the High-Field-Side High-Density (HFSHD) phenomenon observed in other devices [39].
The evolution of this front proceeds in stages. Early in the discharge ( ms), before the density ramp, the radiation is concentrated at the inner divertor target, characteristic of a standard attached divertor [Fig. 7(a)]. As plasma density increases ( ms), an intense radiation zone emerges near the X-point, signaling the birth of the radiative front [Fig. 7(b)]. This event coincides with the onset of the divergent turbulence seen in Fig. 6. With a further increase in density ( ms), this front detaches from the divertor and propagates upward along the HFS separatrix [Fig. 7(c)]. This movement is strongly correlated with the localized edge cooling and turbulence saturation. Finally ( ms), the front moves further inboard and its emissivity intensifies [Fig. 7(d)].
This sequence provides evidence of a developing radiation front on the HFS. While the specific radial location of the peak emissivity has some uncertainty inherent in the tomographic reconstruction, the overall structure and HFS localization of the radiation front are robust features of the data. The role of this dynamic radiative structure will be discussed in detail later, as it appears to be a key triggering event, coinciding with the saturation of the edge density. Notably, the eventual discharge termination was caused by a drop in NBI power that triggered an locked mode and was not a direct consequence of the radiation front.
@@ III.3.2 Radiative Cooling and the Onset of HFS Turbulence
The formation of the HFS radiation front actively cools the plasma and triggers a significant shift in the local turbulent state. Figure 8 provides a quantitative analysis of this transition, using data compiled from multiple discharges. As the HFS radiated power increases beyond a critical threshold, the edge plasma parameters and HFS turbulence changes significantly. The RMS amplitude of the line-integrated HFS density fluctuations (), measured by the CO2 interferometer passing through the radiation front, serves as a proxy for the mode’s strength.
The effect of edge cooling can be characterized by the dimensionless adiabaticity parameter [40], , which represents the comparison between parallel electron motion and turbulent timescales. Here, is the parallel wavenumber of turbulence, is the electron thermal velocity, is the turbulence frequency which can be calculated using the power-weighted mean frequency of the underlying fluctuations, and is the electron-ion collision frequency. When , the plasma enters a non-adiabatic regime where drift-wave instabilities are expected to grow. As shown in Fig. 8(a), the onset of the strong HFS radiation drives to drop sharply below 1.
This transition into a non-adiabatic state coincides with a surge in HFS turbulence. Figure 8(b) shows that the RMS amplitude of HFS density fluctuations () increases sharply as the radiated power crosses the same threshold. The entire event is correlated with a degradation in plasma performance, as indicated by the drop in the confinement factor (color scale) in time sequence. This finding indicates that as the HFS radiation front intensifies, it alters the local plasma state by driving it into a non-adiabatic regime, which is correlated with a surge in turbulence that may ultimately saturate the edge density.